OpenStudy (anonymous):

the number 38.8 is what percent of 97 25 is 2 % of what number

OpenStudy (anonymous):

97 / 38.8 = 2.5 25 / 2 = 2.5 x 100 = 250

OpenStudy (anonymous):

The answer is 250

OpenStudy (anonymous):

its not the answer

OpenStudy (anonymous):

for the 1st one the answer its 40

OpenStudy (gw2011):

97x=38.8 x=38.8/97 x=0.40=40% 0.02x=25 x=25/0.02 x=1250

OpenStudy (anonymous):

n for the 2nd question the answer its 1250